The molar mass of (NH₄)₂CO₃ is ?( Mm N=14 g/mol, H=1 g/mol, C=12 g/mol, O=16 g/mol)


1
Expert's answer
2021-04-28T02:44:41-0400

"N=14"

"H=1"

"C=12"

"O=16"

Molar mass"=[14+(1\times 4)]\times 2=36"

"[12+(16\times 3)]=60"

"36+60=96g\/mol"